# Graph-Based Fault Diagnosis for Rotating Machinery: Adaptive Segmentation and Structural Feature Integration

### v1

@@ 4.1 Datasets
Table 2: Specifications of the datasets used in the analysis.
@@
The dataset(CWRU) bearing dataset used for fault prediction consists of vibration signals collected from bearings under different operating conditions, including different types of faults (ball fault, Inner Race fault, Outer Race fault, and healthy condition) and fault severity (0.007, 0.014 and 0.021 inches). The data were recorded at a sampling frequency of 12 or 48 kHz, meaning each second contained 12,$0$, or 48,000 samples. We use four datasets based on the motor load of 0 Hp, 1 HP, 2 HP, and 3 HP, with approximate data points of 4.8 million each. Table 2 details the dataset used in the analysis.
Table 3: ANOVA and Tukey HSD Results for Vibration Signal RMS Across Load Conditions in the CWRU Dataset
@@
ANOVA All Loads (0 HP, 1 HP, 2 HP, 3 HP) - $F=3966122.36$, $p<0.001$ Yes
Tukey HSD 0 HP vs. 1 HP 0.0495 $<0.001$ [0.0494, 0.0496] Yes
0 HP vs. 2 HP 0.0993 $<0.001$ [0.0992, 0.0995] Yes
0 HP vs. 3 HP 0.1492 $<0.001$ [0.1491, 0.1494] Yes
1 HP vs. 2 HP 0.0498 $<0.001$ [0.0497, 0.0500] Yes
1 HP vs. 3 HP 0.0998 $<0.001$ [0.0996, 0.0999] Yes
2 HP vs. 3 HP 0.0499 $<0.001$ [0.0498, 0.0500] Yes
@@
Table 4: Bearing and Gearbox Fault Types Description (Southeast University DDS Dataset)
Location Type Description
Gearbox Chipped Crack occurs in the gear feet
Miss Missing one of the feet in the gear
Root Crack occurs in the root of gear feet
Surface Wear occurs on the surface of the gear
Bearing Ball Crack occurs in the ball
Inner Crack occurs in the inner ring
Outer Crack occurs in the outer ring
Combination Crack occurs in both the inner and outer rings
Table 5: Statistical Analysis of RMS Vibration Signals for Bearing Conditions (SEU Dataset)
@@
ANOVA (20 Hz–0 V) All Conditions (Ball, Inner, Outer, Health) $F=212.99$ $<0.001$ Yes
Ball vs. Health 0.1098 $<0.001$ [0.0964, 0.1231] Yes
@@
Ball vs. Outer 0.0316 $<0.001$ [0.0183, 0.0449] Yes
Health vs. Inner -0.1171 $<0.001$ [-0.1304, -0.1037] Yes
Health vs. Outer -0.0782 $<0.001$ [-0.0915, -0.0648] Yes
Inner vs. Outer 0.0389 $<0.001$ [0.0256, 0.0522] Yes
T-tests (20 Hz–0 V vs. 30 Hz–2 V) Ball $t=-14.49$ $<0.001$ Yes
Combination $t=5.88$ $<0.001$ Yes
Health $t=4.98$ $<0.001$ Yes
Inner $t=29.32$ $<0.001$ Yes
Outer $t=101.86$ $<0.001$ Yes
@@
An ANOVA test for Ball, Inner, Outer, and Health conditions at 20 Hz–0 V yielded a significant $(F=212.99,p<0.001)$, indicating that bearing condition impacts vibration intensity. Tukey HSD post-hoc tests confirmed significant pairwise differences $(p<0.001)$ for most pairs—Ball vs. Health (mean diff $=0.1098$), Ball vs. Outer (0.0316), Health vs. Inner (-0.1171), Health vs. Outer (-0.0782), and Inner vs. Outer (0.0389)-except Ball vs. Inner (p $=$ 0.494), suggesting similar vibration patterns for these faults.
@@
We perform significance tests on each dataset to ensure the reliability of classification results, especially for simple models like logistic regression used in fault detection. These tests help distinguish real patterns from random noise. A model might seem effective on a dataset, but without statistical validation, its performance could be coincidental. Significance tests provide p-values to confirm that differences—such as RMS variations across fault types—are not due to chance. This validation is crucial in bearing fault diagnosis, where misclassification can cause equipment failure.
@@ 4.2 Experimental Set-up
We trained a logistic regression model on 70% of each dataset and tested it on the remaining 30%, using a detailed evaluation matrix comprising accuracy, precision, recall, and F1-score per fault class. Accuracy reflected the overall correctness of predictions. Precision emphasized the reliability of fault predictions, crucial for reducing false positives in maintenance decisions. Recall ensured actual faults—especially critical types like Inner or Outer—were not missed. F1-score balanced precision and recall, which is essential in imbalanced datasets where Healthy samples often dominate.
To ensure transparency and interpretability, we analyzed the logistic regression coefficients, revealing the influence of graph metrics and node features on predictions—offering actionable insights for engineers. We also visualized modularity, spectral gaps, and computation times to confirm the method’s scalability.
@@ 4.3.1 Optimal Window and Step Size calculation
Figure 2: Optimal window and step size for CWRU dataset with 0 Hp
@@ 4.3.2 Graph Construction
Figure 3: k-Nearest Neighbors (kNN) graph for a cluster split for CWRU 1 Hp Motor Load.
Table 6: Performance Comparison and Graph Properties for CWRU and Gearbox Datasets
@@
Larger subgraphs slightly reduced accuracy and increased runtime, with $k_{\text{max}}=3$, $|\mathcal{G}_{i}|\leq 400$ offering the best small-world alignment. Table 6 summarizes these results. For the SU gearbox dataset, all configurations yielded identical accuracy (1.000); thus, the setup with the lowest construction time was selected.
4.3.3 Classification Analysis and Transfer Learning
The heatmaps in Figure 4(a) illustrate Precision, Recall, and F1-score for 10 fault classes across four load conditions (0 HP, 1 HP, 2 HP, and 3 HP) in the CWRU dataset, using 5-fold cross-validation. Most classes consistently achieve near-perfect scores ($\geq$ 0.999), with minor drops for specific ball faults under 0 HP and 1 HP—e.g., Ball-0.014” shows Recall of 0.923 at 0 HP. The 1 HP load exhibits the most variability, while 3 HP yields the highest overall accuracy (0.997), indicating stronger model performance at higher loads.
(a) Precision, Recall, and F1-score heatmaps for each class across 0 HP, 1 HP, 2 HP, and 3 HP CWRU datasets, evaluated using 5-fold cross-validation.
(b) Classification metric heatmaps (Precision, Recall, and F1-score) across fault types in the combined Gearbox and Bearing datasets. The model maintains consistently high performance with minimal variation.
Figure 4: Heatmaps of class-wise performance metrics for CWRU and Gearbox datasets. Subfigure (a) presents the evaluation on CWRU data under different load conditions. Subfigure (b) shows the results on the combined SU Gearbox dataset .
Figure 4(b) shows classification metrics across combined gearbox and bearing of SU datasets. Each fault type maintains near-perfect Precision, Recall, and F1-score, with slight deviations (e.g., 0.999) in a few fault-load combinations. These results confirm the model’s robustness and generalization capability. Additionally, the model exhibits perfect mean and zero variance across all folds, highlighting its stability and consistency.
(a) Confusion matrix for the CWRU dataset at 0 Hp load.
(b) Confusion matrix for the CWRU dataset at 1 Hp load.
(c) Confusion matrix for the CWRU dataset at 2 Hp load.
(d) Confusion matrix for the CWRU dataset at 3 Hp load.
(e) Confusion matrix for the SU bearing dataset.
(f) Confusion matrix for the SU gearbox dataset.
Figure 5: Confusion matrices depicting the classification performance
The confusion matrices Figure 5 presented illustrate the detailed classification performance of the model on the CWRU dataset under varying load conditions, as well as on the SU bearing and gearbox datasets.
At 0 Hp load (Figure 5(a)), the model exhibits good classification performance but struggles with misclassifications in classes like “Ball-0.007” and “Ball-0.014” due to less distinct vibration signal features, leading to confusion among similar fault types.
Performance markedly improves at 1 Hp load (Figure 5(b)), with nearly all fault types correctly classified and minimal confusion. The confusion matrix’s diagonal dominance reflects enhanced accuracy, likely due to clearer fault signatures from more pronounced vibration features.
At 2 Hp load (Figure 5(c)), the model peaks, achieving near-perfect classification across all classes with virtually no misclassifications. Fault features are most distinguishable at this load, optimizing model input conditions.
Performance remains strong at 3 Hp load (Figure 5(d)) but slightly declines compared to 2 Hp. Minor misclassifications emerge, possibly due to increased system noise or complexity at higher loads, though the confusion matrix still indicates high accuracy across most classes.
In the case of the SU-bearing dataset (Figure 5(e)), the model accurately classifies multiple conditions—including rail faults, combination faults, and healthy states—across varying frequencies and voltage conditions, with confusion matrices showing almost entirely correct predictions. Similarly, in the SU gearbox dataset (Figure 5(f)), the model accurately distinguishes between complex fault types such as chipped tooth, missing tooth, root fault, and surface fault under different operating conditions. The consistency and near-perfect diagonal entries in the confusion matrices for both SU datasets indicate exceptional generalization capability and fault discriminative power.
Overall, the model consistently demonstrates robust and accurate fault diagnosis across various load conditions, effectively distinguishing between closely related fault types even under diverse operating environments. Although the model performs well on individual datasets, real-world fault detection systems often encounter significantly varying operating conditions. Models trained under specific conditions (e.g., a particular machine load) may struggle to generalize across different scenarios. Therefore, transfer learning becomes essential to enable effective fault diagnosis in cross-load scenarios.
Table 7: Transfer Learning Performance of the Proposed Model Across Different Source and Target Configurations
Source Target Accuracy Precision Recall F1-Score
CWRU Bearing Transfer
0 HP Load 0 HP Load 1.000 1.000 1.000 1.000
1 HP Load 0.961 0.961 0.962 0.961
2 HP Load 0.958 0.958 0.959 0.958
3 HP Load 0.948 0.948 0.948 0.948
1 HP Load 0 HP Load 0.987 0.987 0.988 0.987
1 HP Load 0.999 0.999 0.999 0.999
2 HP Load 0.997 0.997 0.997 0.997
3 HP Load 0.986 0.986 0.986 0.986
2 HP Load 0 HP Load 0.974 0.973 0.976 0.974
1 HP Load 1.000 1.000 1.000 1.000
2 HP Load 0.988 0.988 0.988 0.988
3 HP Load 0.980 0.980 0.982 0.980
3 HP Load 0 HP Load 0.983 0.983 0.983 0.983
1 HP Load 0.977 0.978 0.978 0.977
2 HP Load 0.999 0.999 0.999 0.999
3 HP Load 0.999 0.999 0.999 0.999
SU Dataset Transfer
Bearing Gear data Bearing Gear data 1.000a 1.000 1.000 1.000
Bearing Gear data Gearbox Gear data 0.949b 0.949 0.949 0.949
Gearbox Gear data Gearbox Gear data 1.000a 1.000 1.000 1.000
Gearbox Gear data Bearing Gear data 0.973c 0.973 0.973 0.973
a Not significantly different from each other (p ¿ 0.05).
b Significantly different from all others (p ¡ 0.05).
c Significantly different from all others (p ¡ 0.05).
Transfer learning performance using a logistic regression model is evaluated across various source and target load combinations (0 HP, 1 HP, 2 HP, and 3 HP), with metrics including Accuracy, Precision, Recall, and F1-Score. The model demonstrates robust performance across all settings, excelling when source and target loads align. In same-load scenarios—0 HP to 0 HP, 1 HP to 1 HP, 2 HP to 2 HP, and 3 HP to 3 HP—the model achieves perfect classification scores (Accuracy, Precision, Recall, and F1-Score all equal 1.00), underscoring its effectiveness under consistent operating conditions, as shown in Table 7.
In cross-load scenarios, performance remains high but shows slight variability. For instance, transferring from 0 HP to 1 HP yields an Accuracy of 0.961, Precision of 0.961, Recall of 0.962, and F1-Score of 0.961, while 3 HP to 0 HP produces metrics of 0.983 across Accuracy, Precision, Recall, and F1-Score. The lowest performance occurs when transferring from 0 HP to 3 HP, with a Recall, Precision, Accuracy, and F1-Score of 0.948, reflecting challenges in generalizing to significantly different load conditions. Conversely, some cross-load transfers excel, such as 1 HP to 2 HP, achieving an Accuracy and F1-Score of 0.997, and 3 HP to 3 HP, nearing perfect scores of 0.999 across all metrics.
Statistical significance testing reinforces these findings. A paired t-test comparing same-load and cross-load F1-Scores reveals a significant difference ($t=4.56$, $p<0.05$), confirming that same-load scenarios (e.g., 0 HP to 0 HP, F1 = 1.00) outperform the cross-load average (e.g., 0 HP to other loads, mean F1 $\approx$ 0.956). ANOVA tests indicate significant variation in cross-load F1-Scores for source loads at 1 HP ($F\approx 4.12$, $p<0.05$) and 2 HP ($F\approx 3.45$, $p<0.05$). Post-hoc Tukey HSD comparisons show that transfers like 1 HP to 2 HP (F1 = 0.997) significantly outperform 1 HP to 3 HP (F1 = 0.986), and 2 HP to 1 HP (F1 = 0.988) surpasses 2 HP to 0 HP (F1 = 0.974). For source loads at 0 HP and 3 HP, no significant differences emerge among cross-load targets ($p\approx 0.08$ and $p\approx 0.06$, respectively). An overall ANOVA across all cross-load combinations confirms significant differences ($F\approx 5.67$, $p<0.01$), with the largest disparity between 1 HP to 2 HP (F1 = 0.997) and 0 HP to 3 HP (F1 = 0.948), which is statistically significant ($p<0.01$).
Table 7 also details the model’s transfer learning performance, assessed via k-fold cross-validation. The model achieves perfect same-domain performance (1.0 for both Bearing Gear $\to$ Bearing Gear and Gearbox Gear $\to$ Gearbox Gear), ensuring high reliability within known domains. Its strong cross-domain performance, notably Gearbox Gear $\to$ Bearing Gear (Accuracy 0.972806), supports scalability across related systems.

### v2

@@ 3.1 Datasets
Table 2: Specifications of the datasets used in the analysis. Source: Authors own work.
@@
The dataset(CWRU) bearing dataset used for fault prediction consists of vibration signals collected from bearings under different operating conditions, including different types of faults (ball fault, Inner Race fault, Outer Race fault, and healthy condition) and fault severity (0.007, 0.014 and 0.021 inches). The data were recorded at a sampling frequency of 12 or 48 kHz, meaning each second contained 12,0000, or 48,000 samples. We use four datasets based on the motor load of 0 Hp, 1 HP, 2 HP, and 3 HP, with approximate data points of 4.8 million each. Table 2 details the dataset used in the analysis.
Table 3: ANOVA and Tukey HSD Results for Vibration Signal RMS Across Load Conditions in the CWRU Dataset. Source: Authors own work.
@@
ANOVA All Loads (0 HP, 1 HP, 2 HP, 3 HP) - $F=3966122.36$, $p<0.001$ - Yes
Tukey HSD 0 HP vs. 1 HP 0.0495 <0.001<0.001 [0.0494, 0.0496] Yes
0 HP vs. 2 HP 0.0993 <0.001<0.001 [0.0992, 0.0995] Yes
0 HP vs. 3 HP 0.1492 <0.001<0.001 [0.1491, 0.1494] Yes
1 HP vs. 2 HP 0.0498 <0.001<0.001 [0.0497, 0.0500] Yes
1 HP vs. 3 HP 0.0998 <0.001<0.001 [0.0996, 0.0999] Yes
2 HP vs. 3 HP 0.0499 <0.001<0.001 [0.0498, 0.0500] Yes
@@
Table 4: Bearing and Gearbox Fault Types Description (Southeast University DDS Dataset), where each point (*) for each type has an equal number of data points for 20 Hz -0V and 30 Hz - 2V configurations, i.e., 1,048,560 data points for each configuration per type.
Location Type Description Total Data Points
Gearbox Chipped Crack occurs in the gear feet 2,097,120*
Miss Missing one of the feet in the gear 2,097,120*
Root Crack occurs in the root of gear feet 2,097,120*
Surface Wear occurs on the surface of the gear 2,097,120*
Bearing Ball Crack occurs in the ball 2,097,120*
Inner Crack occurs in the inner ring 2,097,120*
Outer Crack occurs in the outer ring 2,097,120*
Combination Crack occurs in both the inner and outer rings 2,097,120*
Table 5: Statistical Analysis of RMS Vibration Signals for Bearing Conditions (SEU Dataset). Source: Authors own work.
@@
ANOVA (20 Hz–0 V) All Conditions (Ball, Inner, Outer, Health) $F=212.99$ <0.001<0.001 - Yes
Ball vs. Health 0.1098 <0.001<0.001 [0.0964, 0.1231] Yes
@@
Ball vs. Outer 0.0316 <0.001<0.001 [0.0183, 0.0449] Yes
Health vs. Inner -0.1171 <0.001<0.001 [-0.1304, -0.1037] Yes
Health vs. Outer -0.0782 <0.001<0.001 [-0.0915, -0.0648] Yes
Inner vs. Outer 0.0389 <0.001<0.001 [0.0256, 0.0522] Yes
T-tests (20 Hz–0 V vs. 30 Hz–2 V) Ball t=14.49t=-14.49 <0.001<0.001 - Yes
Combination $t=5.88$ <0.001<0.001 - Yes
Health $t=4.98$ <0.001<0.001 - Yes
Inner $t=29.32$ <0.001<0.001 - Yes
Outer $t=101.86$ <0.001<0.001 - Yes
@@
An ANOVA test for Ball, Inner, Outer, and Health conditions at 20 Hz–0 V yielded a significant $(F=212.99,p<0.001)$, indicating that bearing condition impacts vibration intensity. Tukey HSD post-hoc tests confirmed significant pairwise differences $(p<0.001)$ for most pairs—Ball vs. Health (mean diff =0.1098=0.1098), Ball vs. Outer (0.0316), Health vs. Inner (-0.1171), Health vs. Outer (-0.0782), and Inner vs. Outer (0.0389)-except Ball vs. Inner (p == 0.494), suggesting similar vibration patterns for these faults.
@@
We perform significance tests on each dataset to ensure the reliability of classification results, especially for simple models used in fault detection. These tests help distinguish real patterns from random noise. A model might seem effective on a dataset, but without statistical validation, its performance could be coincidental. Significance tests provide p-values to confirm that differences—such as RMS variations across fault types—are not due to chance. This validation is crucial in bearing fault diagnosis, where misclassification can cause equipment failure.
@@ 3.2 Experimental Set-up
We employ logistic regression (LR), random forest (RF), and support vector machine (SVM) classifiers to evaluate the effectiveness of the proposed features. Each model is trained on 70% of the data and tested on the remaining 30%. To assess classification performance, we use a comprehensive set of evaluation metrics, including accuracy, precision, recall, and F1-score, computed per fault class. Accuracy measures the overall correctness of predictions, while precision reflects the reliability of fault detection—critical for minimizing false positives in maintenance decisions. Recall captures the model’s ability to correctly identify true fault instances, particularly severe types such as inner and outer defects. The F1-score balances precision and recall, which is especially relevant for imbalanced datasets where healthy samples are more prevalent.
To ensure interpretability, we apply SHAP analysis to quantify the contribution of graph-theoretic metrics and node-level features to model predictions. This provides actionable insights into which structural or statistical characteristics influence classification decisions. Additionally, we visualize confusion matrices, SHAP value distributions, and per-stage computation times to demonstrate the framework’s scalability and practical applicability.
@@ 3.3.1 Optimal Window and Step Size calculation
Figure 2: Optimal window and step size for CWRU dataset with 0 Hp. Source: Authors own work.
@@ 3.3.2 Graph Construction
Figure 3: k-Nearest Neighbors (kNN) graph for a cluster split for CWRU 1 Hp Motor Load. Source: Authors own work.
Table 6: Performance Comparison and Graph Properties for CWRU and Gearbox Datasets. Source: Authors own work.
@@
Larger subgraphs slightly reduce classification accuracy and increase computational cost. The configuration with $k_{\text{max}}=3$ and subgraph size $|\mathcal{G}_{i}|\leq 400$ achieves the best trade-off, aligning well with small-world network characteristics. Table 6 summarizes the performance results using a simple logistic regression (LR) baseline. All configurations achieve perfect accuracy for the SU gearbox dataset (1.000); therefore, the setup with the lowest graph construction time is selected for efficiency.
3.3.3 Classification Analysis
We evaluated the performance of Support Vector Machine (SVM), Logistic Regression (LR), and Random Forest (RF) models on the CWRU and SU bearing fault datasets. Table 7 presents their performance on the CWRU dataset across four horsepower loads (0, 1, 2, and 3 Hp) and SU dataset across gearbox and bearing, comparing accuracy, cross-validation (CV) mean accuracy with standard deviation and inference time.
Table 7: Performance comparison of SVM, Logistic Regression, and Random Forest across horsepower loads for CWRU and SU datasets.
Dataset Model Acc. CV Mean Acc. ($\pm$ SD) Infer. Time (s)
0 Hp SVM 98.7 $98.8\pm 0.9$ 0.0274
Logistic Regression 98.7 $98.8\pm 0.9$ 0.0007
Random Forest 100.0 100.0 $\pm$ 0.000 0.0092
1 Hp SVM 99.0 $99.6\pm 0.1$ 0.2816
Logistic Regression 99.4 $99.6\pm 0.2$ 0.0009
Random Forest 99.9 99.9 $\pm$ 0.1 0.0170
2 Hp SVM 99.3 $99.3\pm 0.2$ 1.3235
Logistic Regression 99.5 $99.6\pm 0.3$ 0.0142
Random Forest 99.7 99.8 $\pm$ 0.2 0.0325
3 Hp SVM 99.8 $99.7\pm 0.2$ 0.7218
Logistic Regression 99.9 $99.7\pm 0.2$ 0.0167
Random Forest 100.0 100.0 $\pm$ 0.1 0.0233
Gearbox SVM 100.0 $100.0\pm 0.000$ 0.001165
Logistic Regression 100.0 $100.0\pm 0.000$ 0.000002
Random Forest 100.0 100.0 $\pm$ 0.000 0.000010
Bearing SVM 99.9 $99.9\pm 0.000$ 0.001074
Logistic Regression 100.0 $100.0\pm 0.000$ 0.000002
Random Forest 100.0 100.0 $\pm$ 0.000 0.000021
The results in Table 7 highlight the strengths and trade-offs of each model for the CWRU dataset. Random Forest consistently achieved the highest accuracy (99.7-100.0) and CV mean accuracy (99.8-100.0, $\pm$0.000-0.2), with perfect classification at 0 and 3 Hps. Logistic Regression delivered strong accuracy (98.7-099.9) and CV mean accuracy (98.8-99.7, $\pm$0.2-0.9) and had the fastest inference times (0.0007-0.0167 s). This speed is critical for real-time applications, such as bearing fault diagnosis, in which rapid decision-making is essential. SVM, while reliable, showed slightly lower accuracy (0.987-0.998) and CV mean accuracy (98.8-99.7, $\pm$0.1-0.9) and the slowest inference times (0.0274-1.3235 s), especially at 2 Hp, making it less suitable for time-sensitive tasks.
Figure 4: Confusion matrices of the Random Forest classifier evaluated on the CWRU bearing dataset under four different motor loads: 0 Hp, 1 Hp, 2 Hp, and 3 Hp. Each matrix illustrates the model’s performance in classifying various bearing fault types (Ball, Inner Race, Outer Race) with different defect sizes, as well as the healthy condition. The strong diagonal dominance across all loads indicates the model’s high accuracy and robustness across varying operating conditions.
Figure 5: Confusion matrices of the Random Forest classifier applied to the SU Bearing dataset.
Figure 6: Confusion matrices of the Random Forest classifier applied to the SU Gearbox dataset
All models exhibited exceptional performance on the SU dataset, which includes the Gearbox and Bearing subsets, reflecting the dataset characteristics, such as well-separated fault classes or less noisy data. For Gearbox, all three models, SVM, Logistic Regression, and Random Forest, achieved perfect accuracy (100.0%) and CV means accuracy (100.0%, ±0.000), indicating robust classification across gearbox-related faults. Logistic Regression stands out with the fastest inference time (0.000002 s), followed closely by Random Forest (0.000010 s), whereas SVM is slower (0.001165 s). For Bearing, Logistic Regression and Random Forest again achieved perfect accuracy (100.0%) and CV mean accuracy (100.0%, ±0.000), with Logistic Regression maintaining its speed advantage (0.000002 s) over Random Forest (0.000021 s) and SVM (0.001074 s). SVM, while nearly perfect (99.9% accuracy and CV mean accuracy), was slightly less accurate on Bearing-SU and consistently slower across both SU subsets.
Random Forest is the most robust and accurate model across the CWRU and SU datasets, making it a top choice for fault diagnosis. On the CWRU dataset, RF consistently achieved the highest accuracy (99.7-100.0%) and cross-validation (CV) mean accuracy (99.8-100.0%, ±0.000-0.2), with perfect classification at 0 and 3 Hp. This demonstrates its exceptional ability to handle complex fault patterns and varying load conditions, ensuring reliable performance in diverse operational scenarios. Similarly, on the SU dataset, RF delivered perfect accuracy (100.0%) and CV mean accuracy (100.0%, ±0.000) for both Gearbox-SU and Bearing-SU, showcasing its capability to excel in datasets with well-separated fault classes. Although its inference times are not the fastest, they remain highly competitive (0.0092-0.0325 s for CWRU, 0.000010–0.000021 s for SU), striking an effective balance between accuracy and computational efficiency.
In contrast, Logistic Regression offers impressive speed, with inference times as low as 0.0007-0.0167 s on CWRU and 0.000002 s on SU, alongside strong accuracy (98.7-100.0% on CWRU, 100.0% on SU). However, its accuracy occasionally falls short of RF, particularly at lower horsepower loads on the CWRU (e.g., 98.7% at 0 Hp), and its performance is less consistent across varying conditions. SVM, while reliable with accuracies of 98.7-99.8% on CWRU and 99.9-100.0% on SU, is significantly hindered by slower inference times (0.0274-1.3235 s on CWRU, 0.001074-0.001165 s on SU), making it less practical for real-time applications. Additionally, SVM’s slightly lower accuracy of SVM on CWRU (e.g., 98.7% at 0 Hp) and Bearing-SU (99.9%) underscores the advantage of RF in precision.
Table 8: Cross-load transfer performance of Random Forest (RF) classifier across different source and target load conditions on the CWRU dataset, and selected scenarios from the SU dataset. Metrics include Accuracy, F1-score, Precision, and Recall.
Source Load Target Load Accuracy F1-score Precision Recall
CWRU Cross-load Transfer
0 Hp 0 Hp 1.000000 1.000000 1.000000 1.000000
1 Hp 0.967655 0.967440 0.968459 0.967655
2 Hp 0.941808 0.941617 0.942108 0.941808
3 Hp 0.963277 0.962806 0.962601 0.963277
1 Hp 0 Hp 0.995614 0.995614 0.995805 0.995614
1 Hp 0.999102 0.999101 0.999109 0.999102
2 Hp 0.995480 0.995475 0.995492 0.995480
3 Hp 0.988701 0.988704 0.988869 0.988701
2 Hp 0 Hp 0.986842 0.986803 0.988421 0.986842
1 Hp 0.997305 0.997309 0.997320 0.997305
2 Hp 1.000000 1.000000 1.000000 1.000000
3 Hp 0.997175 0.997178 0.997188 0.997175
3 Hp 0 Hp 0.982456 0.982375 0.984157 0.982456
1 Hp 0.982929 0.982913 0.982951 0.982929
2 Hp 0.981921 0.981913 0.982003 0.981921
3 Hp 0.999435 0.999435 0.999438 0.999435
SU Cross-load Transfer
Bearing Bearing 1.000000 1.000000 1.000000 1.000000
Gearbox 0.973000 0.911000 0.953100 0.893000
Gearbox Bearing 0.996000 0.857000 0.866000 0.834000
Gearbox 1.000000 1.000000 1.000000 0.990000
To evaluate the Random Forest (RF) model’s effectiveness, we visualized its classification performance using confusion matrices for the CWRU and SU datasets. Figure 4 shows RF’s performance on the CWRU dataset across four motor loads (0–3 Hp) and ten bearing conditions (e.g., Normal, Ball Fault, Inner Race Fault). Figures 5 and 6 highlight RF’s accuracy in identifying fault classes under varying conditions for SU’s bearing and gearbox systems.
3.3.4 Cross Learning
Real-world fault detection often involves diverse operating conditions, where models trained on specific loads may struggle to generalize. Transfer learning is thus critical for effective cross-load fault diagnosis.
Table 9: Cross-Fault Transfer Results: Classification performance for models trained on fault size $0.007\text{\,}\mathrm{{}^{\prime\prime}}$ and tested on $0.007\text{\,}\mathrm{{}^{\prime\prime}}$ and $0.021\text{\,}\mathrm{{}^{\prime\prime}}$ (CWRU dataset).
Source Target Model Accuracy F1-Score Precision Recall
CWRU Cross-Fault Transfer
[…]
RF 1.0000 1.0000 1.0000 1.0000
SVM 0.6667 0.5556 0.5000 0.6667
$007\text{\,}\mathrm{{}^{\prime\prime}} 0.0$$021\text{\,}\mathrm{{}^{\prime\prime}}$ LR 0.1940 0.2446 0.3422 0.1940
RF 0.9718 0.9727 0.9738 0.9718
SVM 0.3446 0.1909 0.3137 0.3446
Figure 7: t-SNE visualization of raw features for the cross-fault transfer scenario from fault size $0.007\text{\,}\mathrm{{}^{\prime\prime}}$ (source) to $0.021\text{\,}\mathrm{{}^{\prime\prime}}$ (target) on the CWRU dataset. Fault types and sizes are represented with distinct colors; markers distinguish source ($\cdot$) and target ($\times$) domains.
In cross-load scenarios, RF demonstrates robust performance on the CWRU dataset, achieving perfect classification (accuracy, F1-score, precision, recall = 1.0) when trained and tested on the same load (e.g., 0 Hp $\to$ 0 Hp). Even when tested on different loads (e.g., 0 Hp $\to$ 1 Hp), RF maintains high performance, with accuracy and F1-scores above 94% (Table 8). For SU’s cross-domain transfers (e.g., bearing $\to$ gearbox), RF achieves up to 99.6% accuracy, with slight F1-score reductions due to domain shifts, underscoring its adaptability across diverse conditions.
The RF model also excels in cross-fault scenarios. On the CWRU dataset, when trained on $0.007\text{\,}\mathrm{{}^{\prime\prime}}$ faults and tested on $0.021\text{\,}\mathrm{{}^{\prime\prime}}$ faults, RF achieves 97.18% accuracy and 97.27% F1-score, significantly outperforming Logistic Regression (LR,  19%) and Support Vector Machine (SVM,  34%) (Table 9). Similarly, on the SU dataset, RF delivers 99.4% accuracy for bearing-to-gearbox transfers and 96.2% for gearbox-to-bearing transfers, surpassing LR ($\leq$82.3%) and SVM ($\leq$90.1%) (Table 10).
Table 10: Cross-Fault Load Classification Performance (SU dataset).
Method Train Domain Test Domain Accuracy F1-Score
RF Bearing (20Hz, 0V) Gearbox (20Hz, 0V) $0.994$ $0.992$
Gearbox (20Hz, 0V) Bearing (20Hz, 0V) $0.962$ $0.955$
SVM Bearing (20Hz, 0V) Gearbox (20Hz, 0V) $0.901$ $0.847$
Gearbox (20Hz, 0V) Bearing (20Hz, 0V) $0.000$ $0.000$
LR Bearing (20Hz, 0V) Gearbox (20Hz, 0V) $0.823$ $0.784$
Gearbox (20Hz, 0V) Bearing (20Hz, 0V) $0.019$ $0.012$
[…]
t-SNE visualizations further illustrate RF’s strengths (Figures 7, 8). These plots show RF’s ability to separate fault types across different sizes and align source and target domains, reflecting its robust feature learning. Unlike LR’s linear constraints or SVM’s sensitivity to margin-based boundaries, RF’s ensemble approach effectively captures complex, nonlinear patterns. Collectively, RF’s high accuracy, robust generalization across loads and fault sizes, and effective domain adaptation (Tables 8, 9, 10; Figures 7, 8) make it an ideal choice for real-world fault diagnosis under varying conditions.